# Nonlinear Schrödinger Network

### v2

@@ Abstract
Deep neural networks (DNNs) have achieved exceptional performance across various fields by learning complex nonlinear mappings from large-scale datasets. However, they encounter challenges such as high computational costs and limited interpretability. To address these issues, hybrid approaches that integrate physics with AI are gaining interest. This paper introduces a novel physics-based AI model called the “Nonlinear Schrödinger Network”, which treats the Nonlinear Schrödinger Equation (NLSE) as a general-purpose trainable model for learning complex patterns including nonlinear mappings and memory effects from data. Existing physics-informed machine learning methods use neural networks to approximate the solutions of partial differential equations (PDEs). In contrast, our approach directly treats the PDE as a trainable model to obtain general nonlinear mappings that would otherwise require neural networks. As a type of physics-AI symbiosis, it offers a more interpretable and parameter-efficient alternative to traditional black-box neural networks, achieving comparable or better accuracy in some time series classification tasks while significantly reducing the number of required parameters. Notably, the trained Nonlinear Schrödinger Network is interpretable, with all parameters having physical meanings as properties of a virtual physical system that transforms the data to a more separable space. This interpretability allows for insight into the underlying dynamics of the data transformation process. Applications to time series forecasting have also been explored. While our current implementation utilizes the NLSE, the proposed method of using physics equations as trainable models to learn nonlinear mappings from data is not limited to the NLSE and may be extended to other master equations of physics.
@@ 1 Introduction
Deep neural networks (DNNs) have emerged as a powerful tool for learning complex nonlinear mappings from data, achieving state-of-the-art performance in various fields including computer vision, natural language processing, and scientific computing. By leveraging their deep hierarchical structure along with billions or even trillions of trainable parameters, DNNs have demonstrated remarkable capability to learn complex patterns from vast amounts of data. However, the success of DNNs comes with challenges, such as the reliance on large-scale datasets, high computational and memory costs, and limited interpretability. To address these challenges, there is a growing interest in hybrid approaches that integrate physics with AI to enable better inference, prediction, vision, control and more [1, 2]. For example, physical principles can be leveraged as prior knowledge to design neural networks with special architectures and loss functions, enabling more efficient modeling of complex physical systems [3, 4]. Moreover, physical systems can act as analog computers to perform specialized computations or transform the data in a manner that reduces the burden on digital processors [5, 6, 7].
One example from our own work is the nonlinear Schrödinger kernel [7], a hardware acceleration technique that was originally proposed to bridge the latency gap between ultrafast data acquisition in optical systems and neural network inference on digital computers. The nonlinear Schrödinger kernel operates by modulating the input data onto the spectrum of a femtosecond laser pulse and then propagating the pulse through a nonlinear optical medium, where complex nonlinear interactions governed by the Nonlinear Schrödinger Equation (NLSE) lead to a transformation of the data that makes it more separable by a simple linear classifier. Notably, operating in the optical domain reduces latency by orders of magnitude compared to utilizing numerical kernels like radius basis function (RBF). The benefits are gained without an increase of data dimensionality. Our follow-up work has demonstrated that while the nonlinear optical system cannot be directly tuned, the nonlinear transformation can be effectively engineered by varying the phase of the input data via spectral phase encoding with an optimization feedback loop [8]. This further improves the performance and generalization of the nonlinear Schrödinger kernel.
While the nonlinear Schrödinger kernel technique has shown promising results in terms of classification accuracy and speed, its physical implementation presents certain limitations, such as the fixed nature of a given optical medium and the need for specialized hardware. However, its underlying principles, particularly the use of nonlinear transformation stemming from the Nonlinear Schrödinger Equation (NLSE) to improve data separability, can potentially be adapted to fully digital algorithms. In this paper, we propose the “Nonlinear Schrödinger Network”, a novel physics-inspired trainable model that operates entirely in the digital domain. This approach treats the NLSE as a trainable model to learn the complex patterns from data. By optimizing the coefficients of the NLSE using backpropagation and stochastic gradient descent (SGD), nonlinear mappings can be learned to transform the data adaptively for classification tasks. Unlike existing work on using neural networks to approximate the solution of a partial differential equation (PDE) [3, 4], our proposed approach directly learns the coefficients of the PDE itself to obtain general nonlinear mappings that would otherwise require neural networks. This physics-inspired approach provides a more interpretable and parameter-efficient model compared to black-box neural networks. By leveraging the low-dimensional nature of physical laws, the Nonlinear Schrödinger Network represents complex data transformations with significantly fewer parameters while achieving accuracy comparable to traditional neural networks on nonlinear classification tasks.
2 Related Work
2.1 Nonlinear Schrödinger Kernel Computing
Despite the increasing speed of digital processors, the execution time of AI algorithms are still orders of magnitude slower than the time scales in ultrafast optical imaging, sensing, and metrology. To address this problem, a new concept in hardware acceleration of AI that exploits femtosecond pulses for both data acquisition and computing called nonlinear Schrödinger kernel has been demonstrated [7]. This method modulates data onto a supercontinuum laser spectrum, which then undergoes a nonlinear optical transformation analogous to a kernel operation, enhancing data classification accuracy. It is shown that the nonlinear optical kernel can improve the linear classification results similar to a traditional numerical kernel (such as the radial-basis-function (RBF)) but with orders of magnitude lower latency. The inference latency of the RBF kernel with a linear classifier is on the order of $10^{-2}$ to $10^{-3}$ second, while the nonlinear Schrödinger kernel achieves a substantially reduced latency on the order of $10^{-5}$ second and better classification accuracy. In the original work [7], the performance is data-dependent due to the limited degrees of freedom and the unsupervised nature of the optical kernel. The follow-up work has demonstrated that by modulating the spectral phase of the input data, the nonlinear optical interactions within the kernel can be effectively engineered [8]. Incorporating this phase encoding scheme within a digital optimization feedback loop allows the optical kernel to be trained in a data-driven manner, minimizing the classification error of the digital backend.
2.2 Neural Networks as PDE Solvers
@@
An emerging trend in blending physics and AI is the use of neural networks as computationally efficient surrogates for solving partial differential equations (PDEs). This approach has the potential to significantly accelerate forward simulations and inverse design used in the optimization of engineered physical systems. For example, neural networks can compute the transmission spectra of metasurfaces orders of magnitude faster than numerically solving the full Maxwell’s equations [9, 10]. Similarly, neural networks trained on real-world data can model intricate fluid flows for applications like weather forecasting and aerodynamics, circumventing the need to directly solve the Navier-Stokes equations [11]. More ambitious attempts have been made to develop general neural network architectures that can learn different types of PDEs. Physics-informed neural network (PINN) [3] is a pioneering example which utilizes a physics-driven loss derived from the equation itself to solve one instance of the PDE with a fixed set of parameters and conditions. Fourier Neural Operator (FNO) [4] takes a step further by learning the nonlinear mapping of an entire family of PDEs.
3 Nonlinear Schrödinger Network
The Nonlinear Schrödinger Equation (NLSE) is one of the canonical equations of physics, accounting for three primary effects: attenuation, dispersion, and nonlinearity. These effects, in the limit of constant group delay dispersion (GDD), are represented by parameters $\alpha$, $\beta_{2}$, and $\gamma$, respectively, as illustrated in Equation 1. The NLSE has demonstrated significant success in describing systems with complex memory and nonlinear effects. In this paper, we treat the NLSE as a trainable model by dividing a virtual medium it into layers, wherein each layer has learnable nonlinear and dispersive properties that can be trained with backpropagation and gradient descent.
@@ 3 Nonlinear Schrödinger Network
Figure 1 presents a schematic diagram of the Nonlinear Schrödinger Network, a novel physics-inspired trainable model. The input data $x$ is treated as an input $E_{in}(t)$ that propagates through a virtual medium, wherein the propagation over a short distance is represented by a Nonlinear Schrödinger Layer with two main components: a linear transformation $D^{(i)}$ parameterized by $\alpha^{(i)}$ and $\beta_{2}^{(i)}$, followed by a nonlinear transformation parameterized by $\gamma^{(i)}$, as described below.
@@
This structure resembles a single block in a common neural network, where a linear layer is typically followed by a nonlinear activation function. By cascading several Nonlinear Schrödinger Layers, we construct a “Nonlinear Schrödinger Network”, a model capable of learning nonlinear mappings from the input data to the desired output.
In Equation 2 and 3, $E(t)$ is a 1D vector with $\omega$ being its corresponding frequency vector. $\alpha^{(i)}$, $\beta_{2}^{(i)}$, and $\gamma^{(i)}$ are trainable scalars in each layer, and $\delta z$ is a constant scalar. $\mathcal{F}$ and $\mathcal{F}^{-1}$ represent 1D Fourier transform and 1D inverse Fourier transform, respectively. As the input data traverses the layers, it undergoes a series of transformations governed by the NLSE, resulting in the transformed output $E_{out}(t)$. Such an output is then mapped to the predicted label $\hat{y}$, which is compared with the true label $y$ to compute the loss $\mathcal{L}$. With all the operations being differentiable, backpropagation and stochastic gradient descent are utilized to update the trainable parameters $\{\alpha^{(i)},\beta_{2}^{(i)},\gamma^{(i)}\}_{i=1}^{M}$ across all the layers.
@@ 4.1 Time Series Classification
Table 1 presents the accuracy and the number of trainable parameters among four different models across the three datasets. The model architectures are visualized in Figure 2. The baseline is a linear classifier consisting of a single densely connected layer. Multi-Layer Perceptron (MLP) is a feedforward neural network with several densely connected hidden layers and nonlinear activations. The 1D Convolutional Neural Network (CNN) includes a few convolutional layers cascaded with densely connected linear layers. In our proposed Nonlinear Schrödinger Network, we add a max pooling layer after the cascaded Nonlinear Schrödinger Layers to mitigate the dimension increase caused by the zero-padding, and a densely connected linear layer to match the output dimension to the number of classes. All models are trained with the Adam optimizer using the cross-entropy loss. As shown in Figure 2, we divide the model architecture into two parts: the “embedding” layers (red dashed box) that maps the original data into a more linearly separable space, and a linear classifier that makes the final decision on the class to which the transformed data belongs. Note that there is no embedding layer in the baseline model.
Figure 2: Model architectures of (a) baseline, (b) Multi-Layer Perceptron (MLP), (c) Convolutional Neural Network (CNN), and (d) Nonlinear Schrödinger Network. The evolutions of data dimensions are visualized in blue rectangles. The “embedding” layers are highlighted with the red dashed box.
@@
00footnotetext: Baseline: a single layer linear classifier. MLP: Multi-Layer Perceptron. CNN: 1D Convolutional Neural Network. Ours: Nonlinear Schrödinger Network. The number of trainable parameters in the “embedding” layers are highlighted in red. All reported accuracies are averaged over 10 random seeds, with standard deviation in parentheses.
@@
We further compare the Nonlinear Schrödinger Network with conventional neural networks, including Multilayer Perceptrons (MLPs) and 1D Convolutional Neural Networks (CNNs). As shown in Table 1, the Nonlinear Schrödinger Network achieves comparable or better accuracy while maintaining significantly fewer parameters. Notably, the number of trainable parameters in the embedding layers of the Nonlinear Schrödinger Network is orders of magnitude fewer than MLP and CNN, as highlighted in red. These results demonstrate the effectiveness of the proposed Nonlinear Schrödinger Layers in capturing the complex nonlinear dynamics of time series data while maintaining a compact model size. This could be a potential solution to the Memory Wall, which is the gap between the performance of the processor and memory in computer architecture. More importantly, the trained Nonlinear Schrödinger Network is an interpretable model with all the parameters having physical meanings as properties of a virtual physical system. This interpretability is a significant advantage over traditional black-box neural networks, as it allows us to gain deeper insights into the underlying dynamics of the data and the decision-making process of the model.

### v3

# Physical Data Embedding for Memory Efficient AI

@@ Abstract
Deep neural networks (DNNs) have achieved exceptional performance across various fields by learning complex, nonlinear mappings from large-scale datasets. However, they face challenges such as high memory requirements and computational costs with limited interpretability. This paper introduces an approach where master equations of physics are converted into multilayered networks that are trained via backpropagation. The resulting general-purpose model effectively encodes data in the properties of the underlying physical system. In contrast to existing methods wherein a trained neural network is used as a computationally efficient alternative for solving physical equations, our approach directly treats physics equations as trainable models. We demonstrate this physical embedding concept with the Nonlinear Schrödinger Equation (NLSE), which acts as trainable architecture for learning complex patterns including nonlinear mappings and memory effects from data. The network embeds data representation in orders of magnitude fewer parameters than conventional neural networks when tested on time series data. Notably, the trained “Nonlinear Schrödinger Network” is interpretable, with all parameters having physical meanings. This interpretability offers insight into the underlying dynamics of the system that produced the data. The proposed method of replacing traditional DNN feature learning architectures with physical equations is also extended to the Gross-Pitaevskii Equation, demonstrating the broad applicability of the framework to other master equations of physics. Among our results, an ablation study quantifies the relative importance of physical terms such as dispersion, nonlinearity, and potential energy for classification accuracy. We also outline the limitations of this approach as it relates to generalizability.
@@ 1 Introduction
Deep neural networks (DNNs) have emerged as a powerful tool for learning complex, nonlinear mappings from data, achieving state-of-the-art performance in various fields including computer vision, natural language processing, and scientific computing. By leveraging their deep hierarchical structure along with billions or even trillions of trainable parameters, DNNs have demonstrated remarkable capability to learn complex patterns from vast amounts of data. However, the success of DNNs comes with challenges, such as reliance on large-scale datasets, high computational and memory costs, and limited interpretability.
As DNN models grow in size and complexity, the memory wall, a fundamental bottleneck in computing architectures, has become increasingly problematic  [2]. In the case of AI data centers, as model parameter counts have increased from millions to hundreds of billions, model memory requirements have severely outpaced on-chip capacity  [3]. This has lead to the adoption of power-hungry solutions such as memory disaggregation and optically connected memory  [4]. Memory bandwidth rather than computational throughput has become a limiting factor, with direct implications for power consumption  [5].
The economic ramifications of this memory wall extend beyond mere technical inefficiency. Data center operators face escalating cooling requirements, increased infrastructure costs, and diminishing returns on computational investments  [6]. The capital expenditure necessary to scale infrastructure for larger models grows non-linearly with model size, challenging the sustainability of continued parameter scaling as a path to improved model performance  [7].
Additionally, the surge in model size further obscures the inner workings of DNNs, undermining interpretability. It is becoming increasingly difficult to trace decisions back to salient features or transparent intermediate representations  [8, 9], posing significant challenges in domains where explainability is crucial, such as healthcare and finance. Despite ongoing efforts in explainable artificial intelligence, the growing complexity of large-scale models continues to hamper interpretability techniques, emphasizing the need for research into new architectures or methodologies that strike a balance between model capacity and transparency  [10].
1.1 Related Work in Physics-AI Symbiosis
To address these challenges, there is a growing interest in hybrid approaches that integrate physics with AI to enable greater efficiency in language, vision, control and more [11, 12, 13, 14]. In contrast to predominantly high-dimensional empirical methodologies in AI, the remarkable success of physics in explaining nature stems from its reliance on low-dimensional, deterministic models that precisely characterize a wide array of phenomena. The design principles that underlie these physical equations can offer strategies for creating more parameter-efficient, interpretable AI approaches. For example, physical principles can be leveraged as prior knowledge to design neural networks with special architectures and loss functions, enabling more efficient modeling  [15, 16, 17]. Moreover, physical systems can act as analog computers to perform specialized computations or transform data in a manner that reduces the burden on digital processors [18, 19, 20, 21]
@@ 1.1 Related Work in Physics-AI Symbiosis
An emerging trend in blending physics and AI is the use of neural networks as computationally efficient surrogates for solving partial differential equations (PDEs). This approach has the potential to significantly accelerate forward simulations and inverse design used in the optimization of engineered physical systems. For example, neural networks can compute the transmission spectra of metasurfaces orders of magnitude faster than numerically solving the full Maxwell’s equations [22, 23]. Similarly, neural networks trained on real-world data can model intricate fluid flows for applications like weather forecasting and aerodynamics, circumventing the need to directly solve the Navier-Stokes equations [24]. More ambitious attempts have been made to develop general neural network architectures that can learn different types of PDEs. The physics-informed neural network (PINN) [15] is an example which utilizes a physics-driven loss derived from the equation itself to solve one instance of the PDE with a fixed set of parameters and conditions. Fourier Neural Operator (FNO) [16] takes a step further by learning the nonlinear mapping of an entire family of PDEs.
2 Physical Embedding with the Nonlinear Schrödinger Equation
Although these approaches hybridize physics and AI, they do not necessary address the core challenges of parameter-efficiency and interpretibility. In contrast, this work proposes the concept of physical embedding, wherein data representations are encoded in the physical properties of a medium rather than distributed across neural network parameters. Unlike existing works which use neural networks to replace PDEs by approximating their solutions  [15, 16], this approach trains a PDE to replace a neural network by directly learning the PDE’s material-dependent coefficients. By leveraging the low-dimensional nature of physical laws, such physical embedding represents complex data with significantly fewer parameters while achieving accuracy comparable to traditional neural networks on nonlinear classification tasks.
In demonstration, we introduce the “Nonlinear Schrödinger Network” (NSN), a novel physics-inspired architecture that operates entirely in the digital domain. This model treats the Nonlinear Schrödinger Equation (NLSE) as a trainable model to learn complex patterns from data. By optimizing the coefficients of the NLSE using backpropagation and stochastic gradient descent (SGD), nonlinear mappings are embedded in the physical parameters of the equation, which transforms the data adaptively for AI tasks.
Figure 1: Comparison of Traditional Convolutional Feature Learning with Physical Embedding. Trainable steps are shown in orange. (a) Convolutional Neural Network with traditional convolutional feature learning layers followed by flattening, nonlinear activation, and linear classifier. (b) Physical Embedding Network wherein a PDE replaces the convolutional layers of a neural network. This embedding utilizes the PDE’s characteristics as the parameters for a deep learning framework. The physical properties of the system are used as learnable parameters, which preserves the elegance and efficiency inherent to the underlying physics. This effectively embeds the complex patterns and relationships of the data into a handful of physically meaningful properties, such as the radius of and distance between gears, in this analogy. The transformed data can then be efficiently separated by a linear classifier.
2.1 Physical Embedding
In traditional neural networks, information is encoded within millions or billions of parameters distributed across a general network architecture  [25]. By composing many simple, trainable operations, these models capture patterns, relationships, and abstractions from the data, essentially embedding representation in the parameters of the network [26].
The concept of physical embedding, introduced here, represents a fundamentally different approach. Rather than learning in an extensive parameter space, the approach trains the physical properties of system governed by a PDE. This effectively embeds the complex patterns and relationships of the data into a handful of physically meaningful properties that describe the system material-dependent dynamics, with the transformation of the data handled by the underlying mechanics of the PDE. As such, the high-dimensional embedding layers of DNNs can be completely replaced by solving the trained PDE for a given input. This process is remarkably parameter efficient, requiring orders of magnitude fewer parameters than conventional neural networks while maintaining comparable performance. Additionally, since the embedding space is constrained by physical laws, the overall learned transformation is interpretable via the original physics through which the embedding equation was designed. Each parameter has a clear physical meaning, offering insight into how the data is being transformed at each layer of the network. One further advantage of this approach is its potential for analog implementation, which can drastically reduce latency compared to traditional digital systems.
In essence, this reframes the problem of representation learning to one of optimizing the coefficients of a predetermined, interpretable physical system. This process bridges the gap between data-driven machine learning and physics-based modeling, offering a more interpretable and parameter-efficient alternative to traditional neural networks. To implement this paradigm in practice, we turn to the Nonlinear Schrödinger Equation (NLSE), a canonical equation in physics that naturally captures a rich variety of nonlinear behaviors and memory effects.
2.2 The Nonlinear Schrödinger Equation as a Neural Network
The Nonlinear Schrödinger Equation (NLSE) is one of the master equations of optics that appears in many areas of mathematical physics and applied science. A nonlinear partial differential equation, it accounts for three primary physical effects: attenuation, dispersion, and nonlinearity. These effects, in the limit of constant group delay dispersion (GDD), are represented by parameters $\alpha$, $\beta_{2}$, and $\gamma$, respectively, as illustrated in Equation 1.
@@ 2.2 The Nonlinear Schrödinger Equation as a Neural Network
Notably, these effects mirror core computational functions in deep neural networks, representing efficient alternatives grounded in fundamental physical processes. Physically, dispersion describes how different spectral components of a wave travel at different velocities, producing pulse reshaping over propagation. From a dynamical systems viewpoint, this introduces a memory effect: the present wave profile depends not only its current amplitude but also on its past configurations. In a neural network analogy, dispersion serves as a proxy for memory–or a form of recurrent context or state–allowing information to mix and influence over space and time. Attenuation captures the loss or dissipation of energy over distance, causing the overall amplitude of the wave to decrease. Numerically, in a neural-network-like framework, this effect parallels a gating mechanism: certain features are scaled down as they travel across layers (propagation steps). Nonlinearity governs how the wave interacts with itself, giving rise to effects such as self-phase modulation or self-focusing. This self-interaction means that the wave’s amplitude can alter its own speed and phase, thereby coupling the wave’s magnitude and phase in a nontrivial way. Computationally, nonlinearity provides the key to representing and learning nonlinear relationships. Like nonlinear activation functions in neural networks, it enables the model to capture richer, more intricate behavior than could be obtained with a purely linear system. Together, these physically grounded mechanisms can be harnessed to learn complex transformations.
@@
Traditionally, the Nonlinear Schrödinger Equation is numerically solved using the split-step Fourier method, which discretizes the propagation medium into alternating linear and nonlinear operators [27]. This solution approach lays the groundwork for the Nonlinear Schrödinger Network, wherein we optimize the parameters of these linear and nonlinear transformations to embed data in a trainable physics-based model. Figure 2 presents a schematic diagram of the Nonlinear Schrödinger Network. The input data $x$ is treated as an input $E_{in}(t)$ that propagates through a virtual medium, wherein the propagation over a short distance is represented by a Nonlinear Schrödinger Layer with two main components: a linear transformation $D^{(i)}$ parameterized by $\alpha^{(i)}$ and $\beta_{2}^{(i)}$, followed by a nonlinear transformation parameterized by $\gamma^{(i)}$, as described below.
@@
This structure resembles a single block in a common neural network, where a linear layer is typically followed by a nonlinear activation function. By cascading several Nonlinear Schrödinger Layers, the Nonlinear Schrödinger Network is capable of learning nonlinear mappings from the input data to the desired output.
In Equation 2 and 5, $E(t)$ is a 1D vector with $\omega$ being its corresponding frequency vector. $\alpha^{(i)}$, $\beta_{2}^{(i)}$, and $\gamma^{(i)}$ are trainable scalars in each layer, and $\delta z$ is a constant scalar. $\mathcal{F}$ and $\mathcal{F}^{-1}$ represent 1D Fourier transform and 1D inverse Fourier transform, respectively. As the input data traverses the layers, it undergoes a series of transformations governed by the NLSE, resulting in the transformed output $E_{out}(t)$. Such an output is then mapped to the predicted label $\hat{y}$, which is compared with the true label $y$ to compute the loss $\mathcal{L}$. With all the operations being differentiable, backpropagation and stochastic gradient descent are utilized to update the trainable parameters $\{\alpha^{(i)},\beta_{2}^{(i)},\gamma^{(i)}\}_{i=1}^{M}$ across all the layers.
@@ 3.1 Time Series Classification
Table 1 presents the accuracy and the number of trainable parameters among four different models across the three datasets. The model architectures are visualized in Figure 3. The baseline is a linear classifier consisting of a single densely connected layer. The Multi-Layer Perceptron (MLP) is a feedforward neural network with several densely connected hidden layers and nonlinear activations. The 1D Convolutional Neural Network (CNN) includes a few convolutional layers cascaded with densely connected linear layers. The Long Short-Term Memory (LSTM) model in Figure 3 is a recurrent neural network that processes input sequences using a 1-layer LSTM. A learned global attention mechanism is used to highlight the most relevant time steps, with the resulting attention scores are normalized via softmax and used to form a single context vector (via a weighted sum of hidden states). After the attention is applied, the output is cascaded through densely connected linear layers. In our proposed Nonlinear Schrödinger Network, we add a max pooling layer after the cascaded Nonlinear Schrödinger Layers to mitigate the dimension increase caused by the zero-padding, and a densely connected linear layer to match the output dimension to the number of classes. All models are trained with the Adam optimizer using the cross-entropy loss. As shown in Figure 3, we divide the model architecture into two parts: the “embedding” layers (red dashed box) that maps the original data into a more linearly separable space, and a linear classifier that makes the final decision on the class to which the transformed data belongs. Note that there is no embedding layer in the baseline model.
Figure 3: Model architectures of (a) baseline (linear classifier), (b) Multi-Layer Perceptron (MLP), (c) Convolutional Neural Network (CNN), (d) Long Short-Term Memory (LSTM) with Attention:.(d) Nonlinear Schrödinger Network. The evolutions of data dimensions are visualized in blue rectangles. The “embedding” layers are highlighted with the red dashed box.
@@
00footnotetext: Baseline: a single layer linear classifier. MLP: Multi-Layer Perceptron. CNN: 1D Convolutional Neural Network. LSTM+Attention: 1-layer Long Short-Term Memory Recurrent Neural Network with Global Attention. Ours: Nonlinear Schrödinger Network. The number of trainable parameters in the “embedding” layers are highlighted in red. All reported accuracies are averaged over 10 random seeds, with standard deviation in parentheses.
@@
We further compare the Nonlinear Schrödinger Network with conventional neural networks, including Multilayer Perceptrons (MLPs), 1D Convolutional Neural Networks (CNNs), and a Long-Short Term Memory (LSTM) recurrent neural network with global attention. As shown in Table 1, the Nonlinear Schrödinger Network achieves comparable or better accuracy while maintaining significantly fewer parameters across time series classification results. The moderate performance of the LSTM model relative to our Nonlinear Schrödinger Network is consistent with observed patterns in time series classification literature. Recurrent Neural Networks (RNNs), including LSTMs, are fundamentally designed for sequential output prediction rather than classification of entire sequences, often struggle with gradient propagation despite their gating mechanisms, and present significant training and parallelization challenges [31]. These architectural limitations likely explain why the LSTM model, despite its substantially larger parameter count, fails to demonstrate consistent performance advantages over our proposed approach across the evaluated datasets.
Notably, the number of trainable parameters in the embedding layers of the Nonlinear Schrödinger Network is orders of magnitude fewer than the other models, as highlighted in red. These results demonstrate the effectiveness of the proposed Nonlinear Schrödinger Layers in capturing the complex nonlinear dynamics of time series data while maintaining a compact model size.